OpenStudy (anonymous):

The cost of coffee is determined by the type of coffee beans that are mixed together. The cost of a local mix of Arabica and Robusta beans is modeled by the equation below, where A = the pounds of Arabica and R = pounds of Robusta: 700A + 1200R = 1,000,000 If a mixture contains 1000 pounds of Arabica beans, how many pounds of Robusta beans should be in the mixture? R = ________________ pounds

OpenStudy (anonymous):

@thatsojessiee please help me

OpenStudy (anonymous):

\[700\times 1,000+1200R=1,000,000\] solve for \(R\)

OpenStudy (anonymous):

i dont get that its the same thing im trying to figure out

OpenStudy (anonymous):

A is arabic beans with a quantity of 1000 which is given

OpenStudy (anonymous):

sorry confusing yes getting it no

OpenStudy (anonymous):

well you got an equation with 2 unknowns R and A right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but in your question A is given

OpenStudy (anonymous):

that is 1000

OpenStudy (anonymous):

ok but r is the confusing part

OpenStudy (anonymous):

R becomes then your only unknown

OpenStudy (anonymous):

so you fill in your A 700*1000 +1200R = 1 000 000

OpenStudy (anonymous):

700 000+ 1200R = 1 000 000

OpenStudy (anonymous):

1200R = 300 000

OpenStudy (anonymous):

R= 300 000/1200

OpenStudy (anonymous):

R=250

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

np